Imagine you wanted to reroute a river. If the restive force of dry ground is 2500 N per square cm of contact and you want the river to run to a half mile away. Assuming that a hand is 18cm in length and 10cm in width, how much work would it take to perform this feat?
Answer #1

Solution) Restive force , F = 2500 N/cm^2

This means for 1 cm^2 force is 2500 N

For Area , A = 18 × 10 cm^2 = 180 cm^2

For 1cm^2 = 2500 N

180 cm^2 = ?

F' = 2500×180 = 450000 N

Displacement S = (1/2) mile

1 mile = 1609.344

(1/2) mile = (1609.344/2) = 804.672 m

S = 804.672 m

Work , W = F.S = (450000).(804.672)

W = 36210.24×10^(4) J

W = 36.21×10^(7) J